Sally travels 400m west 800m south find the displacement and distance?
1
Expert's answer
2020-03-10T11:28:40-0400

The distance


"d=400+800=1200\:\rm m."

The displacement


"s=\sqrt{400^2+800^2}=894\:\rm m."